# SERA: Soft-Verified Efficient Repository Agents

### v1

@@ 5.3 Data Filtering for Specialization
Specializing to Multiple Repositories: We also train on Django and Sympy jointly to investigate whether SERA can be used to specialize to multiple codebases at once. We randomly sample half our initial Django and Sympy datasets from Section 4.3, training on the combined dataset. We jointly evaluate Django and Sympy instances separately, then average them with equal weighting, which debiases against the larger number of Django instances in SWE-Bench Verified. We find that while performance drops slightly on each constituent codebase, the average performance of the combined dataset outperforms $10000$ $\mathbf{T_{1}}$ trajectories from Sera-4.5A-Lite (Table 7). In addition, independent performances on Django and Sympy still compare favorably to the teacher model. This indicates that SERA can be applied to multiple codebases for broadly improved performance, which reflects the needs of enterprises and larger research teams.
@@ 5.4 Teacher Models
Early in Section 4.1, we show that GLM-4.5-Air is a much better teacher than Claude 3.7 Sonnet despite similar SWE-Bench Verified performance. We hypothesize that this is in part due to GLM-4.5-Air’s reasoning traces, which are longer and significantly more elaborate.
To study the effect of reasoning traces, we train on $4200$ $\mathbf{T_{2}}$ trajectories from GLM-4.5-Air where we remove reasoning traces and leave only tool calls. In Table 8 we find that this significantly degrades performance compared to the unchanged trajectories. Our results confirm our hypothesis that high-quality reasoning traces are essential when distilling data for coding agents.
@@ 5.5 Rollout Mixing
Assumptions of ANOVA are also met. A one-way ANOVA reveals no statistically significant difference between trajectory mixing strategies, $F(2,6)=1.91$, $p=.229$, $\eta^{2}=.39$. This indicates that $\mathbf{T_{1}}$ and $\mathbf{T_{2}}$ trajectories can be combined without degrading performance, enabling improved sampling efficiency.
@@ 6 Robustness of Evaluations
Signal-to-Noise Analysis: A practical way to assess whether an observed improvement is real is to compute the signal-to-noise ratio (SNR): the absolute difference between methods divided by the typical run-to-run variance. When SNR <1<1, noise dominates and the result cannot be trusted. When SNR is between 1–2, the result is borderline and requires more seeds. When SNR >2>2, there is likely a real effect. Applying this framework to our experiments:
@@ 7 Deployment
As part of the Ai2 Open Coding Agents release, we provide a lightweight proxy server that enables Claude Code to use SERA as its backend. This section describes the key implementation considerations for deploying SERA in practice.
@@
Infrastructure: The proxy connects Claude Code to any OpenAI-compatible endpoint serving SERA. We use vLLM (Kwon et al., 2023) with the Hermes tool calling parser. For serverless deployment, we provide Modal integration scripts, though the model is portable to any cloud GPU provider or on-premises infrastructure. The proxy handles API translation (Anthropic format to OpenAI format), tool mapping, and response streaming and can be easily modified to handle SERA models specialized on a specific repository. SERA requires at least one 80GB GPU (e.g., A100 80GB, H100, or greater) for deployment. Quantization (e.g., AWQ, GPTQ) can further improve throughput and reduce memory requirements.
@@ 8 Related Work
Training Methods: Beyond data generation, several works explore training approaches for software engineering models. SWE-RL (Wei et al., 2025) applies reinforcement learning on open software evolution data using lightweight rule-based rewards to improve reasoning capabilities. SkyRL-Agent (Cao et al., 2025) provides efficient RL training infrastructure for multi-turn LLM agents. SWE-Fixer (Xie et al., 2025a) combines BM25-based file retrieval with a separate code editing module, training both components on 110K GitHub issues. These reinforcement learning approaches require substantial infrastructure for online rollouts and distributed training, which we discuss further in Section 2.
@@ 9 Limitations
There are several limitations of our work. While we draw certain best-faith conclusions based on our empirical results, this section is mostly about expression reasonable uncertainty and what this might mean for the interpretation of our results. We also highlight the gaps that might make our conclusions one-sided or biased due to the particular experiments that we ran.
@@
Private repo specialization : We demonstrate specialization on Django, Sympy, and Sphinx because these repositories have test data that allows us to evaluate whether specialization works. However, these are public repositories likely included in base model training data. Our specialization experiments may therefore be biased. While specialization effects are well-studied in fine-tuning scaling laws and our results appear plausible, we have not verified specialization on truly private codebases that models have never seen because we have no evaluation data to test this directly.
Statistical robustness concerns of claims: As discussed in Section 6, many of our comparisons are underpowered with $n=3$ seeds. Some reported effects may be noise rather than genuine improvements. We encourage readers to focus on large effects (>3%) and treat smaller differences with appropriate skepticism.
Model-specific results: All experiments use Qwen-3 family of models as the base model and GLM-4.5-Air or GLM-4.6 as teachers. While we have some experiments with Claude 3.7 Sonnet and CLaude 4.0 Sonnet that hints at generalization of our method, we do not know whether our findings generalize to other model families when evaluated thoroughly. The concerns we raise about model-specific optimization in the evaluation section apply equally to our own work.
@@ 10 Broader Impact
We believe that the release of the Ai2 Open Coding Agents that include our SERA models will have significant impact on the research community by enabling research on coing agents without requiring large resources or complicated systems. We also believe private repo specialization will have a significant effect of how small organizations use coding agents. In this section, we extend these discussion and the effects our work has.
Democratizing Coding Agent Research: A central barrier to progress in coding agent research has been the prohibitive cost and infrastructure complexity required to train competitive models. The Ai2 Open Coding Agents initiative aims to remove these barriers, with SERA as its first release. Reinforcement learning approaches require teams of 12 or more researchers, clusters of 64+ GPUs, and months of engineering effort to build sandboxed execution environments. SERA reduces this to a team of five and 16 GPUs, with a total cost under \$2,000. This shift makes coding agent research feasible for individual researchers, small academic labs, and institutions in regions without access to large-scale compute infrastructure. By releasing 200,000 trajectories, our training code, we aim to further lower the barrier to entry so that the study of coding agents is not concentrated among a handful of well-resourced industry labs.
Enabling Private Codebase Specialization:
The ability to specialize a coding agent to a private codebase has significant implications for individuals and small companies. Currently, developers who want AI-assisted coding must send their proprietary code to cloud API providers, creating privacy and intellectual property concerns while these closed system are also not adapted to work well with private data. SERA enables a fundamentally different workflow: users can train a small, local model specialized to their own codebase without exposing their code to any third party. This is particularly relevant for startups with proprietary algorithms, regulated industries (healthcare, finance, defense) where code cannot leave secure environments, and open-source maintainers who want AI assistance tailored to their specific project conventions.
@@ Appendix D Cost Breakdown
Synthetic data generation.
Figure 3 shows scaling curves under three cost regimes: self-hosted inference via vLLM, and API-based inference using GLM-4.5-Air and GLM-4.6 through the z.ai API. To derive the API cost, we analyzed 100 randomly sampled trajectories from the SWE-smith trajectory dataset (Yang et al., 2025) to measure actual token consumption patterns. Each trajectory consists of multiple API calls where the conversation history grows with each turn. For a given API call, the model receives the full conversation history (cached input), the new tool result or observation (uncached input), and produces a response (output). We measured these components and rescaled to 32K context length to match our training setup, yielding an average of 35 API calls per trajectory.
@@
Appendix E Model Card: SERA-32B
This section provides the model card for SERA-32B, following standard practices for documenting machine learning models.
E.1 Model Overview
[…]
E.2 Model Variants
Model HuggingFace Base Teacher SWE-bench Verified
SERA-32B allenai/SERA-32B Qwen 3-32B GLM-4.6 49.5% $\pm$ 1.9%
SERA-32B-GA allenai/SERA-32B-GA Qwen 3-32B GLM-4.5-Air 46.6% $\pm$ 0.7%
SERA-8B allenai/SERA-8B Qwen 3-8B GLM-4.6 31.7% $\pm$ 0.9%
SERA-8B-GA allenai/SERA-8B-GA Qwen 3-8B GLM-4.5-Air 31.7% $\pm$ 0.4%
Table 16: SERA model variants. All results evaluated at 32K context length. Standard deviations computed over 3 random seeds.
E.3 Performance Comparison
Model Type SWE-bench Verified (32K)
SkyRL-8B Open-source 9.4%
Nex-N1-8B Open-source 20.3%
SERA-8B Open-source 31.7%
Qwen 3-32B (base) Open-weight 24.4%
SWE-smith Open-source 32.6%
SkyRL-Agent Open-source 39.4%
DeepSWE Open-source 42.2%
SERA-32B Open-source 49.5%
Devstral-Small-2 (24B) Open-weight 50.0%
GLM-4.5-Air (110B) Open-weight 50.5%
Table 17: Performance comparison on SWE-bench Verified at 32K context. Open-source: code, model weights, and data publicly available. Open-weight: model weights available but training data/code not fully released.
E.4 Quickstart
The easiest way to use SERA is with the sera CLI, which provides seamless integration with Claude Code:
# Install the CLI
uv tool install ai2-sera-cli
# Option 1: Deploy on Modal (recommended for trying out)
modal setup # one-time setup
sera --modal
# Option 2: Use an existing endpoint
export SERA_API_KEY=<your_api_key>
sera --endpoint <endpoint_url>
The first run with --modal takes approximately 10 minutes to download the model ($\sim$65GB) and compile. Subsequent runs start in 1–2 minutes.
E.5 Model Details
Property Value
Developer Allen Institute for AI (Ai2)
Authors Ethan Shen, Daniel Tormoen, Saurabh Shah, Ali Farhadi, Tim Dettmers
Base Model Qwen 3-32B
Teacher Model GLM-4.6 (357B)
Model Type Coding agent / Software engineering agent
Training Method Supervised fine-tuning on synthetic agent trajectories
Context Length 32K tokens
License Apache 2.0
Table 18: SERA-32B model details.
Hyperparameter Value
Epochs 3
Learning Rate 1e-5
Weight Decay 0.01
Max Sequence Length 32,768 tokens
Training Framework Axolotl
Inference Framework vLLM
Compute 40 GPU-days ($\sim$\$2,000)
Table 19: SERA-32B training configuration.
E.6 Training Data
SERA-32B is trained on 200,000 synthetic coding agent trajectories generated using Soft Verified Generation (SVG). SVG is a two-rollout pipeline: (1) a teacher model makes a change to a codebase starting from a randomly selected function, (2) the trajectory is converted into a pull request description, (3) the teacher attempts to reproduce the change given only the PR description, and (4) patches are compared using line-level recall (no test execution required). This approach removes the need for test infrastructure and enables data generation from any repository. The training data was generated from 120 Python codebases using GLM-4.6 (357B) as the teacher model.
E.7 Intended Use
  • Automated software engineering: Bug fixes, feature implementation, refactoring
  • Repository specialization: Fine-tune on private codebases to create specialized coding agents ($\sim$8,000 trajectories / \$1,300)
  • Research: Studying coding agents, data generation methods, and agent behavior
E.8 Limitations
  • SWE-bench training artifact: The model was trained on SWE-bench-style tasks and may attempt to call a nonexistent submit tool when finished editing. The sera-cli proxy handles this automatically.
  • Evaluation scope: Only validated on SWE-bench Verified (Python repositories). Performance on other languages or benchmarks is unknown.
  • Teacher bound: Performance is largely bounded by the teacher model (GLM-4.6) capability.
  • Statistical variance: Results computed over 3 seeds. Effects smaller than 2–3% should be interpreted with caution.
  • Model-specific: Experiments use Qwen 3 as the base model. Generalization to other model families is not validated.
E.9 Bias, Risks, and Limitations
Like any language model without safety filtering, SERA can be prompted to generate harmful or insecure code. Users should be aware of the following risks:
  • Code security: May generate code with security vulnerabilities (e.g., injection attacks, insecure defaults). All generated code should be reviewed before deployment.
  • Accuracy: May produce incorrect or buggy code. Outputs should be tested and verified.
  • Inherited biases: May reflect biases present in the Qwen 3-32B base model and GLM-4.6 teacher model.
  • Misuse potential: Could potentially be used to generate malicious code or identify vulnerabilities for exploitation.
E.10 Responsible Use
This model is intended for research and educational use. Users should adhere to Ai2’s Responsible Use Guidelines (https://allenai.org/responsible-use). Key principles include: use the model for beneficial purposes, review and test all generated code before deployment, do not use to generate malicious software or exploit vulnerabilities, and consider the potential impact of automated code generation in your context.
E.11 Hardware Requirements
SERA-32B requires at least one 80GB GPU (A100 80GB, H100, or equivalent) for deployment at 32K context length. Quantization (AWQ, GPTQ) can reduce memory requirements if needed.
E.12 License
This model is licensed under Apache 2.0. It is intended for research and educational use and may be used commercially in accordance with Ai2’s Responsible Use Guidelines.
E.13 Citation
@article{sera2026,
title={SERA: Soft-Verified Efficient Repository Agents},
author={Shen, Ethan and Tormoen, Daniel and Shah, Saurabh
and Farhadi, Ali and Dettmers, Tim},
year={2026},
institution={Allen Institute for AI},
url={https://allenai.org/papers/opencodingagents}
E.14 Contact
For questions or issues, please contact ethans03@cs.washington.edu or dettmers@cmu.edu, or open an issue at https://github.com/allenai/SERA/issues.

### v2

@@ 5.3 Data Filtering for Specialization
Specializing to Multiple Repositories: We also jointly train on Django and Sympy to investigate whether SERA can be used to specialize to multiple codebases at once. We randomly sample half our initial Django and Sympy datasets from Section 4.3, training on the combined dataset. We evaluate Django and Sympy instances separately, then average them with equal weighting, which debiases against the larger number of Django instances in SWE-Bench Verified. We find that while performance drops slightly on each constituent codebase, the average performance of the combined dataset outperforms 10,000 $\mathbf{T_{1}}$ trajectories from Sera-4.5A-Lite (Table 7). In addition, independent performances on Django and Sympy still compare favorably to the teacher model. This indicates that SERA can be applied to multiple codebases for broadly improved performance, which reflects the needs of enterprises and larger research teams.
@@ 5.4 Teacher Models
In Appendix B, we show that GLM-4.5-Air is a much better teacher than Claude 3.7 Sonnet despite similar SWE-Bench Verified performance. We hypothesize that this is in part due to GLM-4.5-Air’s reasoning traces, which are longer and significantly more elaborate.
To study the effect of reasoning traces, we train on 4,200 $\mathbf{T_{2}}$ trajectories from GLM-4.5-Air where we remove reasoning traces and leave only tool calls. In Table 8 we find that this significantly degrades performance compared to the unchanged trajectories. Our results confirm our hypothesis that high-quality reasoning traces are essential when distilling data for coding agents.
@@ 5.5 Rollout Mixing
We run a one-way ANOVA test revealing no statistically significant difference when trajectories are mixed compared to when they are not: $F(2,6)=1.91$, $p=.229$, $\eta^{2}=.39$. This further indicates that $\mathbf{T_{1}}$ and $\mathbf{T_{2}}$ trajectories can be combined with little to no performance degradation, enabling improved sampling efficiency.
@@ 6 Robustness of Evaluations
Signal-to-Noise Analysis: A practical way to assess the validity of observed improvements is to compute the signal-to-noise ratio (SNR): the absolute difference between methods divided by the typical run-to-run variance. When SNR <1<1, noise dominates and the result cannot be trusted. When SNR is between 1–2, the result is borderline and requires more seeds. When SNR >2>2, there is likely a significant effect. Applying this framework to our experiments:
@@ 7 Deployment
As part of this release, we provide a lightweight proxy server that enables Claude Code to use SERA as its backend. This section describes implementation considerations for deploying SERA in practice.
@@
Infrastructure: The proxy connects Claude Code to any OpenAI-compatible endpoint serving SERA. We use vLLM (Kwon et al., 2023) with the Hermes tool calling parser. For serverless deployment, we provide Modal integration scripts, though the model is portable to any cloud GPU provider or on-premises infrastructure. The proxy handles API translation (Anthropic format to OpenAI format), tool mapping, and response streaming and can be easily modified to handle SERA models specialized on a specific repository. SERA-32B requires at least one 80GB GPU (e.g., A100 80GB, H100, or greater) for deployment. Quantization (e.g., AWQ, GPTQ) can further improve throughput and reduce memory requirements.
@@ 8 Related Work
Training Methods: Beyond data generation, several works explore training approaches for software engineering models. SWE-RL (Wei et al., 2025) applies reinforcement learning on open software evolution data using lightweight rule-based rewards to improve reasoning capabilities. SkyRL-Agent (Cao et al., 2025) provides efficient RL training infrastructure for multi-turn LLM agents. SWE-Fixer (Xie et al., 2025a) combines BM25-based file retrieval with a separate code editing module, training both components on 110K GitHub issues. These reinforcement learning approaches require substantial infrastructure for online rollouts and distributed training, which we discuss in Section 2.
@@ 9 Limitations
There are several limitations of our work. While we draw certain best-faith conclusions based on our empirical results, this section is mostly about uncertainty and what this might mean for the interpretation of our results. We also highlight the gaps that might make our conclusions one-sided or biased due to the particular experiments that we ran.
@@
Private Repository Specialization : We demonstrate specialization on Django, Sympy, and Sphinx because these repositories have test data that allows us to evaluate whether specialization works. However, these are public repositories likely included in base model training data. Our specialization experiments may therefore be biased. While specialization effects are well-studied in fine-tuning scaling laws and our results appear plausible, we have not verified specialization on truly private codebases that models have never seen because we have no evaluation data to test this directly.
Statistical Robustness: As discussed in Section 6, some of our comparisons are underpowered with $n=3$ seeds. Some reported effects may be noise rather than genuine improvements. We encourage readers to focus on large effects (>3%) and treat smaller differences with appropriate skepticism.
Model-Specific Results: All experiments use Qwen-3 family of models as the base model and GLM-4.5-Air or GLM-4.6 as teachers. While we have some experiments with Claude 3.7 Sonnet and Claude 4.0 Sonnet that hint at generalization of our method, we do not know whether our findings generalize to other model families when evaluated thoroughly. The concerns we raise about model-specific optimization in the evaluation section apply equally to our own work.
@@ 10 Broader Impact
We believe that the release of the Ai2 Open Coding Agents that include our SERA models will have significant impact on the research community by enabling research on coing agents without requiring large resources or complicated systems. We also believe private specialization will have a significant effect of how small organizations use coding agents. In this section, we extend these discussion and the effects our work has.
Democratizing Coding Agent Research: A central barrier to progress in coding agent research has been the prohibitive cost and infrastructure complexity required to train competitive models. The Ai2 Open Coding Agents initiative aims to remove these barriers, with SERA as its first release. Reinforcement learning approaches require teams of 12 or more researchers, clusters of 64+ GPUs, and months of engineering effort to build sandboxed execution environments. Our work with SERA presents a much less resource-intensive approach. This shift makes coding agent research feasible for individual researchers, small academic labs, and institutions in regions without access to large-scale compute infrastructure. By releasing 200,000 trajectories, our training code, we aim to further lower the barrier to entry so that the study of coding agents is not concentrated among a handful of well-resourced industry labs.
Enabling Private Codebase Specialization: The ability to specialize a coding agent to a private codebase has significant implications for individuals and small companies. Currently, developers who want AI-assisted coding must send their proprietary code to cloud API providers, creating privacy and intellectual property concerns while these closed system are also not adapted to work well with private data. SERA enables a fundamentally different workflow: users can train a small, local model specialized to their own codebase without exposing their code to any third party. This is particularly relevant for startups with proprietary algorithms, regulated industries (healthcare, finance, defense) where code cannot leave secure environments, and open-source maintainers who want AI assistance tailored to their specific project conventions.
@@ Appendix D Cost Breakdown
Synthetic Data Generation: Figure 3 shows scaling curves under three cost regimes: self-hosted inference via vLLM, and API-based inference using GLM-4.5-Air and GLM-4.6 through the z.ai API. To derive the API cost, we analyzed 100 randomly sampled trajectories from the SWE-smith trajectory dataset (Yang et al., 2025) to measure actual token consumption patterns. Each trajectory consists of multiple API calls where the conversation history grows with each turn. For a given API call, the model receives the full conversation history (cached input), the new tool result or observation (uncached input), and produces a response (output). We measured these components and rescaled to 32K context length to match our training setup, yielding an average of 35 API calls per trajectory.